Imagine that the production function for tuna cans is given by q=6K+ 4L 

where q=Output of tuna cans per hour K=Capital input per hour L= Labor input per hour 


 Assuming capital is fixed at K= 8, how much L is required to produce 100 tuna cans per hour?


1
Expert's answer
2021-07-07T09:06:29-0400

K=8; L=?; q=100

q=6K+ 4L

4L=q-6K

4L=100-6*8

4L=100-48

4L=52

L=52/4

L=13